OpenStudy (anonymous):

Solve for x: x-5/4= 3.

OpenStudy (anonymous):

\[slove for x= \frac{ x-5 }{ 4 }=3\]

OpenStudy (anonymous):

x-5=3*4 hence, x-5=12 so, x=12+5=17